Problem 1 A doctor is checking your hearing in a slightly chilly exam room (T-18 c). a) What is the speed of sound in this exam room? (10pts) b) if the period of the lowest frequency sound you can hear is 0.0005s, what is the frequency of this sound? Has your hearing been affected in some way? How do you know? (15pts) c) What is the wavelength of the sound in this exam room? (10pts)
Answer #1

(a)

Speed of sound in terms of temperature,

v = 331.3 + 0.6*T

At T = 18 oC,

v = 331.3 + 0.6*18

v = 342.1 m/s

(b)

period T = 0.0005 s

Frequency of sound, f = 1 / T = 1 / 0.0005

f = 2000 Hz

(c)

We know that,

wavelength \lambda = v / f

\lambda = 342.1 / 2000

\lambda = 0.171 m